# Unitarity, Causality, and Solar System Bounds, Significantly Limit Using Gravitational Waves to Test General Relativity

### v2

@@ A.4 Fourier Transformations
Here are some Fourier transformations that were useful for finding Newton’s potential from the final amplitude [Donoghue:1994dn]. Most are very well known while others require some $\epsilon$-regularization.
@@ Appendix B Effective action from the Amplitude
We can do this for the rest of the diagrams, namely, contracting the sum of $\mathcal{M}_{\Lambda}$, $\mathcal{M}_{1}$, and $\mathcal{M}_{2}$ with the renormalized cosmological constant $\Lambda_{0}$, with two factors of the graviton propagator yields a non-local Lagrangian as well as the already known counter-Lagrangian. First, the total counter-Lagrangian that corresponds to renormalized $\kappa$ is [Burns:2014bva, tHooft:1974toh]
@@
This allows us to find the following non-local action that matches that of what has been found previously in [Donoghue:2022chi, Donoghue:2014yha], but differs by $3/4$ on the $R^{2}/\Box^{2}$ and $R^{2}\Box$ terms. In fact, any coefficient can be placed on these two terms as long as they have the same value. Having the same value ensures in the limit that $q^{2}\ll m^{2}$ the non-local terms cancel. One thing to note before the final answer is that in contracting over with the two graviton propagators, the non-local function $J(q^{2})$ becomes $B(q^{2})$. The full non-local action is then
@@ Appendix C Discussion on $R^{2}$ and $R_{\mu\nu}R^{\mu\nu}$
In much of the literature when it comes to EFT treatment of gravity, any curvature corrections containing four-derivatives i.e. $R^{2}$, $R_{\mu\nu}R^{\mu\nu}$, and $(\text{Riem})^{2}$ are completely negated. The reasoning for the latter being forgotten is because it can be put into the form of the Gauss-Bonnet term which is a total derivative in the action. The same cannot be said about $R^{2}$ and $R_{\mu\nu}R^{\mu\nu}$. Instead, as in [Stelle:1977ry], if we write down an action of the form $S\supset-\int\sqrt{-g}\left(\alpha R_{\mu\nu}R^{\mu\nu}-\beta R^{2}-\gamma\kappa^{-2}R\right)$, then the coefficients are $\gamma=2$ corresponding to Einstein gravity, while $\alpha$ and $\beta$ can be related to massive bodies as $m_{2}=\sqrt{\gamma/(\alpha\kappa^{2})}$ and $m_{0}=\sqrt{\gamma/(2(3\beta-\alpha)\kappa^{2}}$. If we consider the specific case of a point-particle, then the contribution to Newton’s potential is given by
@@
However, from viewing $R^{2}$ corrections as an EFT, it would appear that $R^{2}$ does not contribute at all as discussed and shown in [AccettulliHuber:2019jqo]. We give a brief derivation of only one of their three different avenues for showing $R^{2}$ does not contribute to Newton’s potential. The argument only requires little group scaling and dimensional analysis.
Consider the 2-to-2 scattering of either scalars or gravitons exchanging a graviton in the background of $R^{2}$. We would expect the amplitude to be factorizable into the sum of two 3-point amplitudes, namely the 2-to-1 scalar to graviton 3-vertex and the all graviton 3-vertex produced by an insertion of $R^{2}$. The latter however can not be modified by $R^{2}$ since all 3-point graviton amplitudes arise from either the expansion of $\sqrt{-g}R$ or $(\text{Riem})^{3}$. As is known from the spinor helicity formalism, all 3-point vertices are uniquely fixed based on the individual particle’s helicity due to little group scaling (for example the general formula can be found in equation (2.99) in [Elvang:2015rqa] [^5] Or for a quick reference, all 3-point polarization amplitudes can be found using $\mathcal{A}_{3}(1^{h_{1}}2^{h_{2}}3^{h_{3}})=c\langle 12\rangle^{h_{3}-h_{1}-h_{2}}\langle 13\rangle^{h_{1}-h_{1}-h_{3}}\langle 23\rangle^{h_{1}-h_{2}-h_{3}}$ ). If we consider the $(1^{+}2^{+}3^{-})$ amplitude (where the rest of single different helicities can be found in a similar manner) we have two possible helicity structures (one found from the other via conjugation) up to a coupling constant
@@
However, our primary focus is on whether or not the inclusion of $\alpha R^{2}$ and $\beta R_{\mu\nu}R^{\mu\nu}$ are valid in any EFT of pure gravity. The couplings $\alpha$ and $\beta$ can in fact run [Donoghue:2014yha], but can still never be found in any scattering experiment. Therefore, based on the arguments above, we would argue that neither $R^{2}$ nor $R_{\mu\nu}R^{\mu\nu}$ contribute to Newton’s potential in any manner when viewing gravitation as an EFT.
[…]

### v3

@@ A.4 Fourier Transformations
Here are some Fourier transformations that were useful for finding Newton’s potential from the final amplitude Donoghue:1994dn . Most are very well known while others require some $\epsilon$-regularization.
@@ Appendix B Effective action from the Amplitude
We can do this for the rest of the diagrams, namely, contracting the sum of $\mathcal{M}_{\Lambda}$, $\mathcal{M}_{1}$, and $\mathcal{M}_{2}$ with the renormalized cosmological constant $\Lambda_{0}$, with two factors of the graviton propagator yields a non-local Lagrangian as well as the already known counter-Lagrangian. First, the total counter-Lagrangian that corresponds to renormalized $\kappa$ is Burns:2014bva ; tHooft:1974toh
@@
This allows us to find the following non-local action that matches that of what has been found previously in Donoghue:2022chi ; Donoghue:2014yha , but differs by $3/4$ on the $R^{2}/\Box^{2}$ and $R^{2}\Box$ terms. In fact, any coefficient can be placed on these two terms as long as they have the same value. Having the same value ensures in the limit that $q^{2}\ll m^{2}$ the non-local terms cancel. One thing to note before the final answer is that in contracting over with the two graviton propagators, the non-local function $J(q^{2})$ becomes $B(q^{2})$. The full non-local action is then
@@ Appendix C Discussion on $R^{2}$ and $R_{\mu\nu}R^{\mu\nu}$
In much of the literature when it comes to EFT treatment of gravity, any curvature corrections containing four-derivatives i.e. $R^{2}$, $R_{\mu\nu}R^{\mu\nu}$, and $(\text{Riem})^{2}$ are completely negated. The reasoning for the latter being forgotten is because it can be put into the form of the Gauss-Bonnet term which is a total derivative in the action. The same cannot be said about $R^{2}$ and $R_{\mu\nu}R^{\mu\nu}$. Instead, as in Stelle:1977ry , if we write down an action of the form $S\supset-\int\sqrt{-g}\left(\alpha R_{\mu\nu}R^{\mu\nu}-\beta R^{2}-\gamma\kappa^{-2}R\right)$, then the coefficients are $\gamma=2$ corresponding to Einstein gravity, while $\alpha$ and $\beta$ can be related to massive bodies as $m_{2}=\sqrt{\gamma/(\alpha\kappa^{2})}$ and $m_{0}=\sqrt{\gamma/(2(3\beta-\alpha)\kappa^{2}}$. If we consider the specific case of a point-particle, then the contribution to Newton’s potential is given by
@@
However, from viewing $R^{2}$ corrections as an EFT, it would appear that $R^{2}$ does not contribute at all as discussed and shown in AccettulliHuber:2019jqo . We give a brief derivation of only one of their three different avenues for showing $R^{2}$ does not contribute to Newton’s potential. The argument only requires little group scaling and dimensional analysis.
Consider the 2-to-2 scattering of either scalars or gravitons exchanging a graviton in the background of $R^{2}$. We would expect the amplitude to be factorizable into the sum of two 3-point amplitudes, namely the 2-to-1 scalar to graviton 3-vertex and the all graviton 3-vertex produced by an insertion of $R^{2}$. The latter however can not be modified by $R^{2}$ since all 3-point graviton amplitudes arise from either the expansion of $\sqrt{-g}R$ or $(\text{Riem})^{3}$. As is known from the spinor helicity formalism, all 3-point vertices are uniquely fixed based on the individual particle’s helicity due to little group scaling (for example the general formula can be found in equation (2.99) in Elvang:2015rqa  [^3] Or for a quick reference, all 3-point polarization amplitudes can be found using $\mathcal{A}_{3}(1^{h_{1}}2^{h_{2}}3^{h_{3}})=c\langle 12\rangle^{h_{3}-h_{1}-h_{2}}\langle 13\rangle^{h_{1}-h_{1}-h_{3}}\langle 23\rangle^{h_{1}-h_{2}-h_{3}}$ ). If we consider the $(1^{+}2^{+}3^{-})$ amplitude (where the rest of single different helicities can be found in a similar manner) we have two possible helicity structures (one found from the other via conjugation) up to a coupling constant
@@
However, our primary focus is on whether or not the inclusion of $\alpha R^{2}$ and $\beta R_{\mu\nu}R^{\mu\nu}$ are valid in any EFT of pure gravity. The couplings $\alpha$ and $\beta$ can in fact run Donoghue:2014yha , but can still never be found in any scattering experiment. Therefore, based on the arguments above, we would argue that neither $R^{2}$ nor $R_{\mu\nu}R^{\mu\nu}$ contribute to Newton’s potential in any manner when viewing gravitation as an EFT.